OpenStudy (anonymous):

If you could help, that would be GREAT! :D Question 1 What is the equation of a vertical line passing through the point (3, –5)? y = –5 x = –5 y = 3 x = 3 Question 2 Solve –2(2x – 3) = –6 x = –3 x = 6 x = 0 x = 3

OpenStudy (anonymous):

2) -4x + 6 = - 6 -4x = - 12 divide both sides by - 4 x = 3

OpenStudy (anonymous):

agreed

OpenStudy (anonymous):

So thats the answer for the seccond one? what about the first?

OpenStudy (anonymous):

Ley y = mx + b

OpenStudy (anonymous):

slope m = ya - yb/xa - xb =.....

OpenStudy (anonymous):

In this case, there is one point!, so

OpenStudy (anonymous):

how do you get rid of the B if i have this problem -5=3x+b

OpenStudy (anonymous):

that isn't right, first you need to find the slope

OpenStudy (anonymous):

ooh... so do the point slope form. ?

OpenStudy (anonymous):

since the slope for a vertical slope is infinite or none, there is no need for m in the equation.....

OpenStudy (anonymous):

so, it becomes y = x + b

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

Yes, I get it now :) thank you!

OpenStudy (anonymous):

just take these notes and remember them, for a vertical line, the slope is infinity or none, for a horizontal line it is 0, for any other line, use the slope formula to figure out m, which is the slope

OpenStudy (anonymous):

good job

OpenStudy (anonymous):

any other problems?

OpenStudy (anonymous):

soo.. I think its x=-5? correct? or is it y=-5?

OpenStudy (anonymous):

and yes i have more problems

OpenStudy (anonymous):

ok then post them and i will help u

OpenStudy (anonymous):

or you can do it this way point slope form y-y1=m(x-x1) where x1 and y1 is your point (x1,y1) we know the point is (3, –5) so x1=3 and y1=-5 and we know that a vertical line has no slope so m=0 plug that in and we have y+5=0(x-3) that right side turns into zero y+5=0 subtract 5 to both sides to get y by itself y=-5 and there you go

OpenStudy (anonymous):

y = -5

OpenStudy (anonymous):

thanks :D

OpenStudy (anonymous):

np

OpenStudy (anonymous):

Question 3 What is the equation of the line passing through (2, –2) and perpendicular to the line y = –x + 3 in slope-intercept form? y = x y = –x y = x – 4 y = x + 4 Question 4 To which graph does the point (–1, –4) belong? y < –x + 4 y < –x – 5 y < 2x – 3 y < 5x + 1 Question 5 What is the standard form equation of the line passing through (5, –3) with a slope of –3? 3x + y = 12 3x + y = –12 3x – y = 12 3x – y = –12

OpenStudy (anonymous):

give me a sec

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

for questions 3 all you have to do is follo my point slope form you have your x1 and y1 and now you need your slope. The slope will be a the negative reciprocal of the slope of the other function given.

OpenStudy (anonymous):

for q 5: let y=mx +c, -3= -15 + c 12=c y= -3x +12, thus 3x + y = 12

OpenStudy (anonymous):

p.s......i think the other guy doesn't know math, but wants to pretend to be smart, lol.....

OpenStudy (anonymous):

for q 4: Y < -X + 4. You put (-1) in for X & (-4) in for Y. (-4) < -(-1) +4. -4 < 1 + 4. -4 < 5 is true. (-4) < -(-1) - 5. -4 < 1 - 5. -4 < -4 is false. -(4) < 2(-1) - 3. -4 < -2 -3. -4 < -5 is false. (-4) < 5(-1) + 1. -4 < -5 + 1. -4 < -4 is false.

OpenStudy (anonymous):

for q 3, i was a bit stuck but now i got it, hold on

OpenStudy (anonymous):

hahaha :P well, i suck at math too :P and thank you for helping mee :) ur the best! lol

OpenStudy (anonymous):

thnxs np, hold on i almost got 3

OpenStudy (anonymous):

First, put the line x+y-1=0 into the slope intercept form

OpenStudy (anonymous):

ex

OpenStudy (anonymous):

your equation is, already in it, my bad

OpenStudy (anonymous):

Next, find the reciprocal of the slope,

OpenStudy (anonymous):

-3, i think

OpenStudy (anonymous):

Next, put into point slope form of the perpendicular line using the point

OpenStudy (anonymous):

y + 2 = 3(x - 2)

OpenStudy (anonymous):

i did this wrong hold on

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

i think its y = x, just by looking at it

OpenStudy (anonymous):

definitely y = x

OpenStudy (anonymous):

any others?

OpenStudy (anonymous):

ok :D and yesss

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

Solve 2 – 5x = x – 10 x = –3 x = –2 x = 2 x = 3 I got x=-2 is that correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oh no

OpenStudy (anonymous):

let me show u

OpenStudy (anonymous):

Question 10 What is the vertex of the graph of y = |–x + 4| + 1? (5, –1) (4, 5) (4, 1) (5, 1) Question 11 Solve 3x + 2y = –12 for y. y = three-halvesx – 6 y = negative three-halvesx – 6 y = three-halvesx – 12 y = negative three-halvesx – 12

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

2 - 5x = x - 10, -5x-x = -10-2, -6x = -12, 6x = 12, x = 2,.....ok?

OpenStudy (anonymous):

oh! okayyyyy :)

OpenStudy (anonymous):

I'm working on 10 and 11 now

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

for q 10: +1 is the y value and, -x+4=0 -x=-4 -x/-1=-4/-4 =4 therefore (4,1)

OpenStudy (anonymous):

ok! thats what i got as well :)

OpenStudy (anonymous):

for q 11; 3x + 2y = –12, 2y = -12 - 3x, y = -3x -12/2,

OpenStudy (anonymous):

well i gtget some sleep, but tell me next time your on open study, and ill help u,k?

OpenStudy (anonymous):

do u know when?

OpenStudy (anonymous):

Okay! sounds great! :D thank you so much for all your help! I really depreciate it :)

OpenStudy (anonymous):

np, message me any more problems k? Im on here every day from 7 - 10 pm ok? ill se yah, p.s. Im from el segundo ca, where u from?

OpenStudy (anonymous):

lol bye

OpenStudy (anonymous):

ok, i will :) and thats cool. haha, im from Florida! lol byee ttyl